Ibsen Company makes two products from a common input. Joint processing costs up to the split-off point total $47,000 a year. The company allocates these costs to the joint products on the basis of their total sales values at the split-off point. Each product may be sold at the split-off point or processed further. Data concerning these products appear below:

Product X Product Y Total
Allocated joint processing costs $ 18,400 $ 28,600 $ 47,000
Sales value at split-off point $ 25,650 $ 37,400 $ 63,050
Costs of further processing $ 22,900 $ 17,200 $ 40,100
Sales value after further processing $ 48,000 $ 55,700 $ 103,700

Required:

a. What is financial advantage (disadvantage) of processing Product X beyond the split-off point? (Negative amount should be indicated by a minus sign.)

b. What is financial advantage (disadvantage) of processing Product Y beyond the split-off point?

c. What is the minimum amount the company should accept for Product X if it is to be sold at the split-off point?

d. What is the minimum amount the company should accept for Product Y if it is to be sold at the split-off point?

Note: List a and b as either financial advantage or financial disadvantage and the values for all (a,b,c,d)

a.
b.
c. Minimum acceptable amount
d. Minimum acceptable amount
Answer #1

a. What is financial advantage (disadvantage) of processing Product X beyond the split-off point?

Sale value after further processing 48000
(-) Sales value at split of point 25650
Incremental revenue 22350
(-) Further processing cost 22900
Disadvantage in further processing -550

What is financial advantage (disadvantage) of processing Product Y beyond the split-off point?

Sale value after further processing 55700
(-) Sales value at split of point 37400
Incremental revenue 18300
(-) Further processing cost 17200
Advantage in further processing 1100

What is the minimum amount the company should accept for Product X if it is to be sold at the split-off point?

= Sale value at split of point - Disadvantage with further processing

= 25650 - 550 = 25100

What is the minimum amount the company should accept for Product Y if it is to be sold at the split-off point?

= Sale value at split of point + Advantage with further processing

37400 + 1100 = 38500

Explanation

In the case of product - Y, we have an advantage with further processing. When this advantage is offered to us at the split off point itself ( i.e 37400 + 1100), then we dont need to go for further processing. Thus 38500 becomes the minimum price for accepting the sale at spilt off point.

In the case of Product - X, we already had best price at split. If it decreases up to 550, then selling at split is not advantagious for us. Thus we determine the minimum price as 25650 - 550